OpenStudy (anonymous):

Write the equation for the area (A) in terms of length (l) of a playground if the width (w) is one half as long as the length (l). You do not need to solve the equation.

OpenStudy (anonymous):

a=l*h they want it in terms of length, l so w = 1/2(length) so A=(l)*[1/2(l)]

OpenStudy (anonymous):

okay I got A=1/2 (1)

OpenStudy (anonymous):

not quite almost you have 1/2 (L) but that is just for the width you also need to multiply by (L) for the length

OpenStudy (anonymous):

A=(L)*1/2(L)

OpenStudy (anonymous):

so then A= 1/2 (1)^2

OpenStudy (anonymous):

that works also

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

no worries